# Turning Up the Heat: Min-p Sampling for Creative and Coherent LLM Outputs

### v2

@@ 7 Conclusion
In this paper, we introduced min- sampling, a novel truncation sampling method for large language models that dynamically adjusts the sampling threshold based on the modelβs confidence at each decoding step. Our approach effectively balances creativity and coherence, particularly at higher temperatures where traditional methods like top- sampling often struggle.
Through comprehensive experiments across diverse benchmarks, we demonstrated that min- sampling consistently outperforms existing methods in both quality and diversity of outputs. Extensive human evaluations further confirmed a strong preference for min- sampling over top-, highlighting its practical advantages in real-world applications.
The key strengths of min- sampling are its simplicity, computational efficiency, and ease of integration into existing pipelines. By enabling models to generate text that is both creative and coherent, min- sampling addresses the longstanding trade-off between diversity and quality in text generation.
Min- sampling represents a significant advancement in generative language modeling, potentially enhancing a wide range of applications requiring high-quality and diverse text generation.
@@
We have made significant efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in AppendixΒ B and is also available anonymously at our project repository.333hhttps://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΒ 5 and SectionΒ B.1. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
@@
We believe the benefits of entropy and uncertainty-based methods outweigh these risks. We strongly encourage safety and alignment research leveraging uncertainty and entropy, as this can significantly benefit robustness, truthfulness and reduced hallucinations (Stolfo etΒ al., 2024; Wang & Zhou, 2024).
@@ Acknowledgements
We thank Apart Research, particularly Esben Kran and Jason Schreiber, for providing extensive guidance throughout our Apart Lab Fellowship, during which min- evolved from an interesting idea to a full, empirically validated research paper. We also thank Wand AI for their generous sponsorship and assistance with evaluations on Mistral Large and Llama3 70B.
@@
AB conceived the original idea of min- sampling, tested and shared it with hobbyists within the open-source LLM community, leading to its rapid and widespread external adoption.
MN initiated, planned and wrote most of the draft and final versions, conducted extensive GPQA and GSM8K evaluations on Mistral 7B, and coordinated with coauthors, collaborators and inference providers throughout the research process.
CN acted as research supervisor under the Apart Research Fellowship, oversaw initial drafts, handled paper formatting, early revisions and data visualization.
AK conducted GSM8K Self-Consistency evaluations to assess Diversity-Accuracy tradeoffs and helped flesh out theoretical explanations of min-.
AR conducted evaluations on Mistral Large and Llama 3 70B, coordinated research collaboration with Wand AI, generated outputs for human evaluation, and provided insights on optimal settings.
RSZ rewrote a substantial portion of the final draft, proposed the inclusion of a human evaluation, and provided guidance on evaluations, data presentation and academic publishing standards.
Appendix A Limitations and Future Work
While min- sampling shows significant promise, there are limitations and opportunities for future research:
Generalization to Other Models:
Our experiments focused on the Mistral 7B and Mistral Large language models. Future work should explore the effectiveness of min- sampling with larger models and different architectures to assess its generalizability.
Hyperparameter Sensitivity:
The base probability threshold is a critical hyperparameter. Investigating methods for dynamically adjusting based on context or developing guidelines for optimal settings across various tasks could enhance performance. This was particularly challenging, as MAUVE hyperparameter sweeps are measured without temperature scaling on GPT2-XL (Pillutla etΒ al., 2023), and require pre-selection on exact hyperparameters. min-, however, is a novel method often used in conjunction with temperature scaling without extensive literature/experimental data. We discuss this further in SectionΒ B.1.
Theoretical Analysis:
A deeper theoretical understanding of why min- sampling performs better, particularly at high temperatures, could provide insights into the behavior of language models and guide the development of even more effective sampling strategies. A possible direction can be using information theory (Shwartz-Ziv etΒ al., 2023)
Applicability to Other Domains:
Research into high-temperature regimes:
High-temperature regimes have been underexplored relative to low-temperature regimes. Min- sampling hopes to unlock exploration, experimentation, and applications in such areas.
Human Evaluation Scope:
Our human evaluation involved participants selecting pre-generated outputs. We note that min-βs popularity within the open source community for creative writing is interactive in nature; hence, we hope for adoption on interactive human evaluation platforms such as the Chatbot Arena (Chiang etΒ al., 2024).
Combining uncertainty and CoT decoding methods:
This mirrors our hypothesis that choosing high-certainty tokens enables more accurate final answers, while diverse token choices benefit intermediate reasoning steps. For example, we note that GPQA scores on Mistral 7B increased from to , but only with Temperature Only and min-. With the recent release of OpenAIβs O1 models, which leverage CoT methods at inference for advanced reasoning capabilities, we note several novel decoding methods that combine uncertainty and CoT sampling approaches to improve model reasoning in a simple manner, with minimal added overhead and architectural changes (Wang & Zhou, 2024; Shrek & Frog, (2024) . We aim to actively explore such methods in future work.
Appendix B Min- Implementation and Documentation
@@
This implementation code, along with other similar implementations in other open-source inference engines, logs of automated evaluations for GPQA, GSM8K Chain-of-Thought and AlpacaEval Creative Writing is available at https://anonymous.4open.science/r/minp_paper-767F/.
B.1 Hyperparameters Settings
@@
For min-, base probability thresholds of and were used, while top- sampling employed .
@@
B.2 Test Time Compute challenges
While running GPQA and GSM8K CoT for and sampling via Hugging Face and the EleutherAI Evaluation Harness (Gao etΒ al., 2023), we noted that test-time compute increased exponentially with temperature. Throughout our experiments, min-, top-, and top- generally took 2-5 minutes to evaluate on Mistral 7B at every temperature for both GPQA and GSM8K. Runtime on an A100 Colab increased from Β 5 minutes at to 8 minutes at and 30 minutes at . Neither nor seemed to function at . On GSM8K, and each took 2 hours to evaluate, which is equal to the average for our Llama 70B/Mistral Large run. This time also appeared to scale exponentially with increased temperature.
@@ B.2 Test Time Compute challenges
B.3 How percentages thresholds differ for min- and top-
@@
Consider the following top 5 tokens probabilities: 80%, 7%, 3%, 2%, 1%. With top- set to 0.9, the top 3 tokens comprising 90% of the distribution is preserved. With min- = 0.1, and the resulting truncation threshold at 8%, only the top token is preserved, and 20% of the original distribution is truncated.
@@
Appendix C Additional Results
C.1 Ablation Study

### v4

@@ 7 Conclusion
In this paper, we introduced min- sampling, a novel truncation method for LLMs that dynamically adjusts sampling thresholds based on model confidence. Our approach effectively balances creativity and coherence, particularly at higher temperatures where methods like Top- sampling struggle.
Through comprehensive experiments across diverse benchmarks, we demonstrate that min- sampling consistently outperforms existing methods in both quality and diversity. Human evaluations confirm these advantages, showing a clear preference for min- outputs in real-world applications.
Min- samplingβs strengths lie in its simplicity, computational efficiency, and seamless integration into existing pipelines. By addressing the longstanding trade-off between diversity and quality, min- represents a significant advancement for generative language modeling, enhancing applications that require both high-quality and diverse text generation. For more detailed examples, see sectionΒ D.1.
@@
We have made significant efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in SectionΒ A.1 and is also available at our project repository.444https://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΒ 5 and SectionΒ A.2. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
Additionally, due to min-βs popularity within the open-source community, we note that there are multiple, unaffiliated and independent replications of the benchmarks and claims presented in this paper. Most notably, our claims of min- improving medium-to-high temperature GPQA and GSM8K evaluation results have been replicated by Top-N sampling developers (Tang etΒ al., 2024), and independent AlpacaEval Creative Writing and EQ-Bench evaluations were conducted by Ilya Gusev (Gusev, 2023; 2024), with results broadly confirming our findings about min-βs advantages.
@@
β’
AI Safety and Interpretability: We are actively exploring follow-up work applying Min- to mechanistic interpretability for AI safety and alignment. Specifically, we are investigating its use in uncertainty-aware generation, neuron activation filtering, and structured latent selection to enhance model robustness and truthfulness. We discuss an extension of this approach in Β§D.2, where we introduce min-, a variation that leverages standard deviation-based filtering for broader applications.
We believe the benefits of entropy and uncertainty-based methods outweigh these risks. We strongly encourage safety and alignment research leveraging uncertainty and entropy, as this can significantly benefit robustness, truthfulness, and reduced hallucinations (Stolfo etΒ al., 2024; Wang & Zhou, 2024).
@@ Acknowledgements
We thank Apart Research, particularly Esben Kran and Jason Schreiber, for providing extensive guidance throughout our Apart Lab Fellowship, during which min- evolved from an interesting idea to a full, empirically validated research paper. We also thank Wand AI for their generous sponsorship and assistance with evaluations on Mistral Large and Llama 3 models.
@@
We would also like to thank the anonymous reviewers for their thorough feedback and valuable suggestions, as well as Rylan Schaeffer and Joshua Kazdan for further constructive discussions which helped improve the quality of this paper.
@@
Andrew Baker conceived the original idea of min- sampling, tested and shared it with hobbyists within the open-source LLM community, leading to its rapid and widespread external adoption.
Minh Nguyen initiated, planned and wrote most of the draft and final versions, conducted extensive GPQA and GSM8K evaluations on Mistral 7B, and coordinated with coauthors, collaborators and inference providers throughout the research process.
Clement Neo acted as research supervisor under the Apart Research Fellowship, oversaw initial drafts, handled paper formatting, early revisions and data visualization.
Andreas Kirsch conducted GSM8K Self-Consistency evaluations to assess Diversity-Accuracy tradeoffs and helped flesh out theoretical explanations of min-.
Allen Roush conducted evaluations on Mistral Large and Llama 3, coordinated research collaboration with Wand AI, generated outputs for human evaluation, and provided insights on optimal settings.
Ravid Shwartz-Ziv rewrote a large portion of the final draft, proposed the inclusion of a human evaluation, and provided guidance on evaluations, data presentation and academic publishing standards.
Appendix A Implementation Details
A.1 Min- Implementation (code listing)
@@
This implementation code, along with other similar implementations in other open-source inference engines, logs of automated evaluations for GPQA, GSM8K Chain-of-Thought and AlpacaEval Creative Writing is available at https://github.com/menhguin/minp_paper/.
A.2 Hyperparameters Settings (Selection Methodology)
@@
For min-, base probability thresholds of and were used, while Top- sampling employed and .
@@
A.3 Test Time Compute challenges
While running GPQA and GSM8K CoT for and sampling via Hugging Face and the EleutherAI Evaluation Harness (Gao etΒ al., 2023), we noted that test-time compute increased exponentially with temperature. During our experiments, min-, Top-, and top- generally took 2-5 minutes to evaluate on Mistral 7B at every temperature for both GPQA and GSM8K. Runtime on an A100 Colab increased from Β 5 minutes at to 8 minutes at and 30 minutes at . Neither nor functioned at . On GSM8K, and each took 2 hours to evaluate, equal to the average for our Llama 70B/Mistral Large run. This time also scaled exponentially with increased temperature.
@@ A.3 Test Time Compute challenges
A.4 How percentages thresholds differ for Min-p and Top-p
@@
Consider the following top 5 token probabilities: 80%, 7%, 3%, 2%, 1%. With Top- set to 0.9, the top 3 tokens comprising 90% of the distribution is preserved. With min- = 0.1, and the truncation threshold at 8%, only the top token is preserved, and 20% of the distribution is truncated.
@@
A.5 Detailed Community Adoption Statistics
To provide further transparency regarding the adoption of Min- sampling, TableΒ 6 summarizes key statistics from major projects that have integrated the method.
| Repository | Stars | Downstream Repos |
| Primary Frameworks | ||
| Hugging Face Transformers | 140,000 | 291,000 |
| Ollama | 130,000 | 400 |
| vLLM | 39,700 | β |
| SGLang | 11,000 | β |
| unsloth | 32,800 | β |
| Additional Implementations | ||
| llama.cpp | 75,600 | β |
| open-webui | 80,000 | β |
| text-generation-webui | 42,700 | β |
| continue | 24,000 | β |
| Total | 575,800 | 291,400+ |
Clarification on Community Adoption Metrics:
We wish to clarify that our earlier claim specifying "54,000 repositories and 1.1 million stars" was based on preliminary GitHub scans using search terms like "min_p" which produced many false positives. An independent contributor has provided a detailed analysis of the limitations of this search approach Nguyen (2025).
For this revised assessment, we focused on identifying verified integrations in high-traffic frameworks where Min- is explicitly implemented as a sampling option in the official codebase. This approach, while potentially undercounting total adoption, provides a more conservative and verifiable measure of Min-βs impact in the open-source ecosystem.
Given the difficulty of obtaining direct usage statistics, as many leading providers do not specifically collect or disclose such data, we believe this evidence of integration into major frameworks represents the most reliable indicator of Min-βs practical adoption and impact.
Appendix B Benchmark Evaluation Results
B.1 Ablation Study